Circular No. 7070

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998ez IN ANONYMOUS GALAXY L. Germany reports the discovery by the Mount Stromlo Abell Cluster

Supernova Search Team (cf. IAUC 6639) of an apparent supernova (V = 20) on V and R CCD images taken on Dec. 8 (limiting mag V = 21) by S. Sabine at the Mount Stromlo 1.27-m telescope (+ Macho Camera). SN 1998ez is located near Abell 3188 at α = 3^{h}57^{m}32^{s}.93, δ = −27^{o}12′13′′.6 (equinox 2000.0), which is 0′′.7 east and 4′′.0 north of the host galaxy’s center. A nearby star has position end figures 30^{s}.25, 11′16′′.1. SN 1998ez was not visible on similar images taken on Oct. 24 (limiting mag V = 20.5) but it was confirmed by S. Sabine with the same telescope (limiting mag V = 20), and by L. Germany with the Mount Stromlo 2.3-m telescope (in B, V, R, and I), on Dec. 11.

COMET P/1998 W2 (HERGENROTHER) The following orbital elements from MPEC 1998-Y15 show this to be

a short-period comet:

T = 1998 Dec. 6.057 TT  ω = 13.^{o}972
e = 0.60389  Ω = 356.555 }2000.0
q = 1.41866 AU  i = 21.915
a = 3.58147 AU n^{o} = 0.145416 P = 6.78 years

SUPERNOVA 1998aq IN NGC 3982 CCD magnitude estimates, unfiltered unless otherwise noted: May 9.94

UT, 13.5 (S. Moretti and S. Tomaselli, Forlí, Italy); 22.863, 13.4 (G. Masi, Ceccano, Italy); Nov. 20.139, V = 19.0 ± 0.2 (L. Kiss and K. Sárneczky, Piszkéstető, Konkoly Observatory, Hungary).

SUPERNOVAE 1998bu, 1998ec, 1998ef, 1998eg, 1998en, 1998es, 1998et

Additional CCD magnitudes by L. Kiss and K. Sárneczky, Piszkéstető, Konkoly Observatory, Hungary (unfiltered unless otherwise noted): SN 1998bu in NGC 3368, Nov. 20.189 UT, V = 17.8 ± 0.1; SN 1998ec in UGC 3576, Nov. 24.137, 18.4 ± 0.2; SN 1998ef in UGC 646, Oct. 26.85, R = 14.95 ± 0.10 (with G. Szabó); SN 1998eg in UGC 12133, Oct. 26.69, R = 17.70 ± 0.05 (with G. Szabó); SN 1998en in UGC 3645, Nov. 24.130, 18.5 ± 0.2; SN 1998es in NGC 632, Nov. 22.698, 14.70 ± 0.05; SN 1998et, Nov. 24.160, 19.0 ± 0.2.

1998 December 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green